# On linearisation and uniqueness of preduals

### v2

@@ 1. Introduction
The purpose of this paper is twofold. First, we present a general mechanism how to transfer a strong linearisation of a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of scalar-valued functions on a non-empty set $\Omega$ to a vector-valued counterpart. Second, we characterise those spaces among the spaces $\mathcal{F}(\Omega)$ which have a (strongly) unique predual in certain classes of locally convex Hausdorff spaces. Strong linearisations are special preduals. Recall that a locally convex Hausdorff space $X$ is called a dual space and a tuple $(Y,\varphi)$ a predual of $X$ if the tuple consists of a locally convex Hausdorff space $Y$ and a topological isomorphism $\varphi\colon X\to Y_{b}^{\prime}$ where $Y_{b}^{\prime}$ is the strong dual of $Y$. In [25] we derived necessary and sufficient conditions for the existence of preduals and strong linearisations of bornological spaces such that the predual has certain properties like being complete and barrelled, a DF-space, a Fréchet space or completely normable (cf. Proposition 3.5 and Corollary 3.6). A strong linearisation of a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ where $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ is a triple $(\delta,Y,T)$ of a locally convex Hausdorff space $Y$ over the field $\mathbb{K}$, a map $\delta\colon\Omega\to Y$ and a topological isomorphism $T\colon\mathcal{F}(\Omega)\to Y_{b}^{\prime}$ a strong linearisation of $\mathcal{F}(\Omega)$ if $T(f)\circ\delta=f$ for all $f\in\mathcal{F}(\Omega)$ (see [10, p. 683], [17, p. 181, 184] and [25, Proposition 2.6, p. 5]).
@@ 2. Notions and preliminaries
The preceding definition of a predual is already given in [25, Definition 2.1, p. 3] and for Banach spaces e.g. in [9, p. 321]. In the setting of Banach spaces the definition of the equivalence of preduals is given in [13, Definition 2.1]. It is also easily checked that $\sim$ actually defines an equivalence relation on the family of all preduals of a dual space. Further, if we have a predual $(Y,\varphi)$ of a dual space $X$ and another locally convex Hausdorff space $Z$ which is topologically isomorphic to $Y$, we can always augment $Z$ to a predual which is equivalent to $(Y,\varphi)$.
@@
Now, if we want to study whether a dual space has a unique predual by identification via topological isomorphisms, we have to restrict the range of preduals we consider because even a Banach space may have a predual which is also a Banach space, and another predual which is not a Banach space (see [25, Example 3.15, p. 11]). Two such preduals cannot be topologically isomorphic.
@@ 3. Linearisation and uniqueness
We recall from [25, p. 21–22] that a topological space $\Omega$ is said to be a $gk_{\mathbb{R}}$-space if for any completely regular space $Y$ and any map $f\colon\Omega\to Y$, whose restriction to each compact $K\subset\Omega$ is continuous, the map is already continuous on $\Omega$. Examples of Hausdorff $gk_{\mathbb{R}}$-spaces are Hausdorff $k$-spaces, metrisable spaces, locally compact Hausdorff spaces and strong duals of Fréchet–Montel spaces (DFM-spaces). Further, for a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ we define $\delta_{x}\colon\mathcal{F}(\Omega)\to\mathbb{K}$, $\delta_{x}(f)\coloneqq f(x)$, for $x\in\Omega$.
3.6 Corollary ([25, Corollary 4.1, Remark 4.3, Theorem 4.10, p. 19, 22]).
@@
We note that the conditions $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for some $\widetilde{\tau}\leq\tau_{\operatorname{p}}$ are also necessary for the existence of a strong complete barrelled linearisation of a bornological function space $\mathcal{F}(\Omega)$ (see [25, Theorem 4.5, p. 19]).
@@
Then it holds (a)normal-$\Rightarrow$(b)normal-$\Leftrightarrow$(c)normal-$\Leftrightarrow$(d). If the family $(\delta(x))_{x\in\Omega}$ is linearly independent, then it holds (b)normal-$\Rightarrow$(a).
@@
and hence $\lambda_{0}(\delta(x))=(\Phi_{\varphi}^{-1}\circ\Phi_{T})(\delta(x))$, implying our claim as $(\Phi_{\varphi}^{-1}\circ\Phi_{T})_{\mid Y_{0}}\colon Y_{0}\to Z_{0}$ is a topological isomorphism by Proposition 2.3. Due to [18, 3.4.2 Theorem, p. 61–62] and the density of $Y_{0}$ in the complete space $Y$ by [25, Proposition 2.7, p. 6] we can uniquely extend $\lambda_{0}$ to a topological isomorphism $\lambda\colon Y\to\overline{Z_{0}}$ where $\overline{Z_{0}}$ denotes the closure of $Z_{0}$ in the complete space $Z$.
@@
Then it holds (a)normal-$\Rightarrow$(b)normal-$\Leftrightarrow$(c)normal-$\Leftrightarrow$(d). If the family $(\delta(x))_{x\in\Omega}$ is linearly independent, then it holds (b)normal-$\Rightarrow$(a).
@@
The condition on $\mathcal{F}(\Omega)$ in Remark 3.11 is for example fulfilled if $\Omega\subset\mathbb{C}$ and $\mathcal{F}(\Omega)$ contains all polynomials of degree less than $|\Omega|-1$ where $|\Omega|$ denotes the cardinality of $\Omega$.
@@ 4. Linearisation of vector-valued functions
Let $\mathcal{F}(\Omega)$ be a locally convex Hausdorff space of $\mathbb{K}$-valued functions on a non-empty set $\Omega$, $Y$ a locally convex Hausdorff space over the field $\mathbb{K}$ and $\delta\colon\Omega\to Y$. Then condition (i) of [25, Proposition 2.5, p. 5] is equivalent to $\delta\in\mathcal{F}(\Omega,Y)_{\sigma}$.
@@
for all $u\in L(Y,E)$ where we used [27, Proposition 22.14, p. 256] for the last equation. Therefore $\chi(u)\in\mathcal{F}(\Omega,E)_{\sigma,b}$, which means that the map $\chi$ is well-defined. The map $\chi$ is also linear and thus the estimate above implies that it is continuous. The map $\chi$ is also injective as the span of $\{\delta(x)\;|\;x\in\Omega\}$ is dense in $Y$ by [25, Proposition 2.7, p. 6].
@@
for every $e^{\prime}\in E^{\prime}$, yielding $\Psi(f)(\delta(x))\in E$ for every $x\in\Omega$. By [25, Proposition 2.7, p. 6] the span of $\{\delta(x)\;|\;x\in\Omega\}$ is dense in $Y$. Thus for every $y\in Y$ there is a net $(y_{\iota})_{\iota\in I}$ in this span such that it converges to $y$ and $\Psi(f)(y_{\iota})\in E$ for each $\iota\in I$. For every $p\in\Gamma_{E}$ we have
@@
Due to weak-strong principles for holomorphic, harmonic and Lipschitz continuous functions special cases of Theorem 4.5 and Remark 4.4 for Banach spaces $E$ over $\mathbb{C}$ were already obtained before, for instance in [29, Theorem 2.1, p. 869] for the Banach space $\mathcal{F}(\Omega)=\mathcal{H}^{\infty}(\Omega)$ of bounded holomorphic functions on an open subset $\Omega$ of a Banach space, in [7, Lemma 10, p. 243] for a Banach space $\mathcal{F}(\mathbb{D})$ of holomorphic functions on the open unit disc $\mathbb{D}\subset\mathbb{C}$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact, in [26, Lemma 5.2, p. 14] for a Banach space $\mathcal{F}(\mathbb{D})$ of harmonic functions on $\mathbb{D}$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact, in [19, Proposition 6, p. 3], which also covers [16, Theorem 3.1 (Linearization Theorem), p. 128] and [1, Proposition 3.1 (c), p. 11], for a closed subspace $\mathcal{F}(\Omega)$ of the Banach space $\mathcal{H}v(\Omega)$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact where $\Omega$ is an open connected subset of a Banach space and $v\colon\Omega\to(0,\infty)$ a continuous function, and in [11, Theorem 1, p. 239–240] where $\mathcal{F}(\Omega)=\mathcal{H}_{b}(\Omega)$ is the space of holomorphic functions defined on a balanced open subset $\Omega$ of a complex normed space which are bounded on $\Omega$-bounded sets. $\mathcal{H}_{b}(\Omega)$ is a Fréchet space if equipped with the topology of uniform convergence on $\Omega$-bounded sets, and we note that its predual $Y\coloneqq\mathfrak{P}_{b}(\Omega)$ in [11, p. 238–239] is an inductive limit of a sequence of Banach spaces and thus ultrabornological, in particular barrelled, by construction. For Banach spaces $E$ over $\mathbb{C}$ Theorem 4.5 and Remark 4.4 also generalise [3, Lemma 2.3.1, p. 67] where $(\delta,Y,T)\coloneqq(\Delta,G,J)$ is a strong Banach linearisation of a Banach space $\mathcal{F}(\Omega)$ of $\mathbb{C}$-valued functions on a non-empty set $\Omega$ with Banach predual $(G,J)$ such that $\delta_{x}\in G$ for all $x\in\Omega$, as well as [3, Corollary 2.3.4, p. 68] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is the weighted Fréchet space of holomorphic functions on a complex Banach space $\Omega$ and $\mathcal{V}\coloneqq(v_{n})_{n\in\mathbb{N}}$ an increasing sequence of continuous functions $v_{n}\colon\Omega\to(0,\infty)$. Further, Theorem 4.5 for Banach spaces $E$ over $\mathbb{C}$ also covers [2, Theorem 2, p. 283] where $\mathcal{F}(\Omega)=\mathcal{VH}(\Omega)$ is the inductive limit of the weighted Banach spaces $\mathcal{H}v_{n}(\Omega)$ of holomorphic functions on a complex Banach space $\Omega$ and $\mathcal{V}\coloneqq(v_{n})_{n\in\mathbb{N}}$ a decreasing sequence of continuous functions $v_{n}\colon\Omega\to(0,\infty)$ if $\mathcal{VH}(\Omega)$ satisfies $(\operatorname{CNC)}$ for $\tau_{\operatorname{co}}$ or $(\mathcal{VH}(\Omega)_{\mathcal{B},\tau_{\operatorname{co}}},\beta)$ is distinguished by [25, Corollary 4.7, p. 20–21]. For complete locally convex Hausdorff spaces $E$ over $\mathbb{C}$ Theorem 4.5 also generalises [5, 3.7 Proposition, p. 292] combined with [5, 1.5 Theorem (e), p. 277–278] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is the weighted space of holomorphic functions on a balanced open subset of $\Omega\subset\mathbb{C}^{d}$ and $\mathcal{V}$ a family of radial upper semicontinuous functions $v\colon\Omega\to[0,\infty)$ such that the weighted topology $\tau_{\mathcal{V}}$ induced by $\mathcal{V}$ fulfils $\tau_{\operatorname{co}}\leq\tau_{\mathcal{V}}$ (see [5, p. 272, 274]), $\mathcal{HV}(\Omega)$ is bornological and the polynomials are contained in $\mathcal{HV}_{0}(\Omega)\coloneqq\{f\in\mathcal{HV}(\Omega)\;|\;\forall\;v\in\mathcal{V}:\;fv\;\text{vanishes at}\;\infty\}$. However, we note that in contrast to Theorem 4.5 quasi-complete $E$ are allowed in [5, 3.7 Proposition, p. 292]. Theorem 4.5 for complete locally convex Hausdorff spaces $E$ over $\mathbb{C}$ also improves [6, Theorem 3.3, p. 35] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is bornological, $\Omega$ an open connected subset of $\mathbb{C}^{d}$ and $\mathcal{V}$ a point-detecting Nachbin family of continuous non-negative functions on $\Omega$. If $E$ is a complete locally convex Hausdorff space and $\mathcal{F}(\Omega)=\mathscr{L}(F\times G,\mathbb{K})$ the space of continuous bilinear forms on the product $\Omega\coloneqq F\times G$ of two locally convex Hausdorff spaces $F$ and $G$, then Theorem 4.5 also covers [15, Chap. I, §1, ${}^{\circ}$1, Proposition 2, p. 30–31] (cf. [18, 15.1.2 Theorem, p. 325]) if the projective tensor product $Y\coloneqq F\otimes_{\pi}G$ is quasi-barrelled (instead of $F\otimes_{\pi}G$ one may also use its completion $F\widehat{\otimes}_{\pi}G$ by [18, 3.4.2 Theorem, p. 61–62]). For instance, $F\otimes_{\pi}G$ is (quasi-)barrelled (and $F\widehat{\otimes}_{\pi}G$ barrelled by [18, 11.3.1 Proposition (e), p. 223]) if $F$ and $G$ are metrisable and barrelled by [18, 15.6.6 Proposition, p. 337], or if $F$ and $G$ are quasi-barrelled DF-spaces by [18, 15.6.8 Proposition, p. 338]. However, we emphasize that [15, Chap. I, §1, ${}^{\circ}$1, Proposition 2, p. 30–31] in contrast to Theorem 4.5 does not need the restrictions that $E$ is complete and $F\otimes_{\pi}G$ quasi-barrelled. Moreover, Theorem 4.5 also complements [10, Theorem 3, p. 690] where a special continuous linearisation $(e,\mathcal{F}_{\ast}(\Omega),L)$ of a linear space $\mathcal{F}(\Omega)$ of scalar-valued continuous functions on a non-empty topological Hausdorff space $\Omega$ instead of a strong linearisation is considered and it is shown that the map $L(\mathcal{F}_{\ast}(\Omega),E)\to\mathcal{F}(\Omega,E)_{\sigma}\cap\mathcal{C}(\Omega,E)$, $u\mapsto u\circ e$, is an algebraic isomorphism if $E$ is a complete locally convex Hausdorff space. Here, $\mathcal{C}(\Omega,E)$ denotes the space of continuous maps from $\Omega$ to $E$. We refer the reader to [17, p. 182–184] for a summary of the construction of $(e,\mathcal{F}_{\ast}(\Omega),L)$
@@
If $Y$ is complete, then $Y$ is also a free object generated by $\Omega$ in the category $\mathcal{C}_{2}$ of complete locally convex Hausdorff spaces with continuous linear operators as morphisms in the sense of [12, p. 2] once the spaces $\Omega$ belong to a category $\mathcal{C}_{1}$ such that $\delta$ and any $f\in\mathcal{F}(\Omega,E)_{\sigma,b}$ is a morphism of $\mathcal{C}_{1}$ and there is a “forgetful” functor from $\mathcal{C}_{2}$ to $\mathcal{C}_{1}$.
@@
Let $(E,\tau)$ be a bornological locally convex Hausdorff space, $\mathcal{B}$ the family of $\tau$-bounded sets, $\widetilde{\tau}$ a locally convex Hausdorff topology on $E$ and $\widetilde{\gamma}$ denote the finest locally convex Hausdorff topology on $E$ which coincides with $\widetilde{\tau}$ on $\tau$-bounded sets. Suppose that a subset of $E$ is $\tau$-bounded if and only if it is $\widetilde{\gamma}$-bounded (see [25, Definition 3.12, p. 10]). Then $(E,\widetilde{\gamma})^{\prime}$ determines boundedness of $(E,\tau)$. Indeed, $(E,\widetilde{\gamma})^{\prime}\subset(E,\tau)^{\prime}$ by [25, Remark 3.13 (c), p. 10–11]. Let $B\subset E$ be $\sigma(E,(E,\widetilde{\gamma})^{\prime})$-bounded. Then $B$ is $\widetilde{\gamma}$-bounded by Mackey’s theorem. It follows that $B$ is $\tau$-bounded by assumption.
@@
Let $(\mathcal{F}(\Omega),\tau)$ be a complete bornological DF-space of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ satisfying $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for some $\widetilde{\tau}$ such that $\tau_{\operatorname{p}}\leq\widetilde{\tau}$, $U\subset\Omega$ a set of uniqueness for $\mathcal{F}(\Omega)$, $E$ a complete locally convex Hausdorff space and $G\subset E^{\prime}$ determine boundedness. If $f\colon U\to E$ is a function such that $e^{\prime}\circ f$ admits an extension $f_{e^{\prime}}\in\mathcal{F}(\Omega,E)_{\sigma}$ for each $e^{\prime}\in G$, then there exists a unique extension $F\in\mathcal{F}(\Omega,E)_{\sigma}$ of $f$.

### v3

@@ 1. Introduction
The purpose of this paper is twofold. First, we present a general mechanism how to transfer a strong linearisation of a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of scalar-valued functions on a non-empty set $\Omega$ to a vector-valued counterpart. Second, we characterise those spaces among the spaces $\mathcal{F}(\Omega)$ which have a (strongly) unique predual in certain classes of locally convex Hausdorff spaces. Strong linearisations are special preduals. Recall that a locally convex Hausdorff space $X$ is called a dual space and a tuple $(Y,\varphi)$ a predual of $X$ if the tuple consists of a locally convex Hausdorff space $Y$ and a topological isomorphism $\varphi\colon X\to Y_{b}^{\prime}$ where $Y_{b}^{\prime}$ is the strong dual of $Y$. In [25] we derived necessary and sufficient conditions for the existence of preduals and strong linearisations of bornological spaces such that the predual has certain properties like being complete and barrelled, a DF-space, a Fréchet space or completely normable (cf. Proposition 3.5 and Corollary 3.6). A strong linearisation of a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ where $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ is a triple $(\delta,Y,T)$ of a locally convex Hausdorff space $Y$ over the field $\mathbb{K}$, a map $\delta\colon\Omega\to Y$ and a topological isomorphism $T\colon\mathcal{F}(\Omega)\to Y_{b}^{\prime}$ if $T(f)\circ\delta=f$ for all $f\in\mathcal{F}(\Omega)$ (see [10, p. 683], [17, p. 181, 184] and [25, Proposition 2.6, p. 1595]).
@@ 2. Notions and preliminaries
The preceding definition of a predual is already given in [25, Definition 2.1, p. 1593] and for Banach spaces e.g. in [9, p. 321]. In the setting of Banach spaces the definition of the equivalence of preduals is given in [13, Definition 2.1]. It is also easily checked that $\sim$ actually defines an equivalence relation on the family of all preduals of a dual space. Further, if we have a predual $(Y,\varphi)$ of a dual space $X$ and another locally convex Hausdorff space $Z$ which is topologically isomorphic to $Y$, we can always augment $Z$ to a predual which is equivalent to $(Y,\varphi)$.
@@
Now, if we want to study whether a dual space has a unique predual by identification via topological isomorphisms, we have to restrict the range of preduals we consider because even a Banach space may have a predual which is also a Banach space, and another predual which is not a Banach space (see [25, Example 3.15, p. 1601]). Two such preduals cannot be topologically isomorphic.
@@ 3. Linearisation and uniqueness
We recall from [25, p. 1611–1612] that a topological space $\Omega$ is said to be a $gk_{\mathbb{R}}$-space if for any completely regular space $Y$ and any map $f\colon\Omega\to Y$, whose restriction to each compact $K\subset\Omega$ is continuous, the map is already continuous on $\Omega$. Examples of Hausdorff $gk_{\mathbb{R}}$-spaces are Hausdorff $k$-spaces, metrisable spaces, locally compact Hausdorff spaces and strong duals of Fréchet–Montel spaces (DFM-spaces). Further, for a locally convex Hausdorff space $\mathcal{F}(\Omega)$ of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ we define $\delta_{x}\colon\mathcal{F}(\Omega)\to\mathbb{K}$, $\delta_{x}(f)\coloneqq f(x)$, for $x\in\Omega$.
3.6 Corollary ([25, Corollary 4.1, Remark 4.3, Theorem 4.10, p. 1609, 1612]).
@@
We note that the conditions $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for some $\widetilde{\tau}\leq\tau_{\operatorname{p}}$ are also necessary for the existence of a strong complete barrelled linearisation of a bornological function space $\mathcal{F}(\Omega)$ (see [25, Theorem 4.5, p. 1609]).
@@
Then it holds (a)$\Rightarrow$(b)$\Leftrightarrow$(c)$\Leftrightarrow$(d). If the family $(\delta(x))_{x\in\Omega}$ is linearly independent, then it holds (b)$\Rightarrow$(a).
@@
and hence $\lambda_{0}(\delta(x))=(\Phi_{\varphi}^{-1}\circ\Phi_{T})(\delta(x))$, implying our claim as $(\Phi_{\varphi}^{-1}\circ\Phi_{T})_{\mid Y_{0}}\colon Y_{0}\to Z_{0}$ is a topological isomorphism by Proposition 2.3. Due to [18, 3.4.2 Theorem, p. 61–62] and the density of $Y_{0}$ in the complete space $Y$ by [25, Proposition 2.7, p. 1596] we can uniquely extend $\lambda_{0}$ to a topological isomorphism $\lambda\colon Y\to\overline{Z_{0}}$ where $\overline{Z_{0}}$ denotes the closure of $Z_{0}$ in the complete space $Z$.
@@
Then it holds (a)$\Rightarrow$(b)$\Leftrightarrow$(c)$\Leftrightarrow$(d). If the family $(\delta(x))_{x\in\Omega}$ is linearly independent, then it holds (b)$\Rightarrow$(a).
@@
The condition on $\mathcal{F}(\Omega)$ in Remark 3.11 is for example fulfilled if $\Omega\subset\mathbb{C}$ and $\mathcal{F}(\Omega)$ contains all polynomials of degree less than $|\Omega|$ where $|\Omega|$ denotes the cardinality of $\Omega$.
@@ 4. Linearisation of vector-valued functions
Let $\mathcal{F}(\Omega)$ be a locally convex Hausdorff space of $\mathbb{K}$-valued functions on a non-empty set $\Omega$, $Y$ a locally convex Hausdorff space over the field $\mathbb{K}$ and $\delta\colon\Omega\to Y$. Then condition (i) of [25, Proposition 2.5, p. 1595] is equivalent to $\delta\in\mathcal{F}(\Omega,Y)_{\sigma}$.
@@
for all $u\in L(Y,E)$ where we used [27, Proposition 22.14, p. 256] for the last equation. Therefore $\chi(u)\in\mathcal{F}(\Omega,E)_{\sigma,b}$, which means that the map $\chi$ is well-defined. The map $\chi$ is also linear and thus the estimate above implies that it is continuous. The map $\chi$ is also injective as the span of $\{\delta(x)\;|\;x\in\Omega\}$ is dense in $Y$ by [25, Proposition 2.7, p. 1596].
@@
for every $e^{\prime}\in E^{\prime}$, yielding $\Psi(f)(\delta(x))\in E$ for every $x\in\Omega$. By [25, Proposition 2.7, p. 1596] the span of $\{\delta(x)\;|\;x\in\Omega\}$ is dense in $Y$. Thus for every $y\in Y$ there is a net $(y_{\iota})_{\iota\in I}$ in this span such that it converges to $y$ and $\Psi(f)(y_{\iota})\in E$ for each $\iota\in I$. For every $p\in\Gamma_{E}$ we have
@@
Due to weak-strong principles for holomorphic, harmonic and Lipschitz continuous functions special cases of Theorem 4.5 and Remark 4.4 for Banach spaces $E$ over $\mathbb{C}$ were already obtained before, for instance in [29, Theorem 2.1, p. 869] for the Banach space $\mathcal{F}(\Omega)=\mathcal{H}^{\infty}(\Omega)$ of bounded holomorphic functions on an open subset $\Omega$ of a Banach space, in [7, Lemma 10, p. 243] for a Banach space $\mathcal{F}(\mathbb{D})$ of holomorphic functions on the open unit disc $\mathbb{D}\subset\mathbb{C}$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact, in [26, Lemma 5.2, p. 14] for a Banach space $\mathcal{F}(\mathbb{D})$ of harmonic functions on $\mathbb{D}$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact, in [19, Proposition 6, p. 3], which also covers [16, Theorem 3.1 (Linearization Theorem), p. 128] and [1, Proposition 2.3 (c), p. 3029], for a closed subspace $\mathcal{F}(\Omega)$ of the Banach space $\mathcal{H}v(\Omega)$ such that its closed unit ball is $\tau_{\operatorname{co}}$-compact where $\Omega$ is an open connected subset of a Banach space and $v\colon\Omega\to(0,\infty)$ a continuous function, and in [11, Theorem 1, p. 239–240] where $\mathcal{F}(\Omega)=\mathcal{H}_{b}(\Omega)$ is the space of holomorphic functions defined on a balanced open subset $\Omega$ of a complex normed space which are bounded on $\Omega$-bounded sets. $\mathcal{H}_{b}(\Omega)$ is a Fréchet space if equipped with the topology of uniform convergence on $\Omega$-bounded sets, and we note that its predual $Y\coloneqq\mathfrak{P}_{b}(\Omega)$ in [11, p. 238–239] is an inductive limit of a sequence of Banach spaces and thus ultrabornological, in particular barrelled, by construction. For Banach spaces $E$ over $\mathbb{C}$ Theorem 4.5 and Remark 4.4 also generalise [3, Lemma 2.3.1, p. 67] where $(\delta,Y,T)\coloneqq(\Delta,G,J)$ is a strong Banach linearisation of a Banach space $\mathcal{F}(\Omega)$ of $\mathbb{C}$-valued functions on a non-empty set $\Omega$ with Banach predual $(G,J)$ such that $\delta_{x}\in G$ for all $x\in\Omega$, as well as [3, Corollary 2.3.4, p. 68] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is the weighted Fréchet space of holomorphic functions on a complex Banach space $\Omega$ and $\mathcal{V}\coloneqq(v_{n})_{n\in\mathbb{N}}$ an increasing sequence of continuous functions $v_{n}\colon\Omega\to(0,\infty)$. Further, Theorem 4.5 for Banach spaces $E$ over $\mathbb{C}$ also covers [2, Theorem 2, p. 283] where $\mathcal{F}(\Omega)=\mathcal{VH}(\Omega)$ is the inductive limit of the weighted Banach spaces $\mathcal{H}v_{n}(\Omega)$ of holomorphic functions on a complex Banach space $\Omega$ and $\mathcal{V}\coloneqq(v_{n})_{n\in\mathbb{N}}$ a decreasing sequence of continuous functions $v_{n}\colon\Omega\to(0,\infty)$ if $\mathcal{VH}(\Omega)$ satisfies $(\operatorname{CNC)}$ for $\tau_{\operatorname{co}}$ or $(\mathcal{VH}(\Omega)_{\mathcal{B},\tau_{\operatorname{co}}},\beta)$ is distinguished by [25, Corollary 4.7, p. 1610–1611]. For complete locally convex Hausdorff spaces $E$ over $\mathbb{C}$ Theorem 4.5 also generalises [5, 3.7 Proposition, p. 292] combined with [5, 1.5 Theorem (e), p. 277–278] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is the weighted space of holomorphic functions on a balanced open subset of $\Omega\subset\mathbb{C}^{d}$ and $\mathcal{V}$ a family of radial upper semicontinuous functions $v\colon\Omega\to[0,\infty)$ such that the weighted topology $\tau_{\mathcal{V}}$ induced by $\mathcal{V}$ fulfils $\tau_{\operatorname{co}}\leq\tau_{\mathcal{V}}$ (see [5, p. 272, 274]), $\mathcal{HV}(\Omega)$ is bornological and the polynomials are contained in $\mathcal{HV}_{0}(\Omega)\coloneqq\{f\in\mathcal{HV}(\Omega)\;|\;\forall\;v\in\mathcal{V}:\;fv\;\text{vanishes at}\;\infty\}$. However, we note that in contrast to Theorem 4.5 quasi-complete $E$ are allowed in [5, 3.7 Proposition, p. 292]. Theorem 4.5 for complete locally convex Hausdorff spaces $E$ over $\mathbb{C}$ also improves [6, Theorem 3.3, p. 35] where $\mathcal{F}(\Omega)=\mathcal{HV}(\Omega)$ is bornological, $\Omega$ an open connected subset of $\mathbb{C}^{d}$ and $\mathcal{V}$ a point-detecting Nachbin family of continuous non-negative functions on $\Omega$. If $E$ is a complete locally convex Hausdorff space and $\mathcal{F}(\Omega)=\mathscr{L}(F\times G,\mathbb{K})$ the space of continuous bilinear forms on the product $\Omega\coloneqq F\times G$ of two locally convex Hausdorff spaces $F$ and $G$, then Theorem 4.5 also covers [15, Chap. I, §1, n1, Proposition 2, p. 30–31] (cf. [18, 15.1.2 Theorem, p. 325]) if the projective tensor product $Y\coloneqq F\otimes_{\pi}G$ is quasi-barrelled (instead of $F\otimes_{\pi}G$ one may also use its completion $F\widehat{\otimes}_{\pi}G$ by [18, 3.4.2 Theorem, p. 61–62]). For instance, $F\otimes_{\pi}G$ is (quasi-)barrelled (and $F\widehat{\otimes}_{\pi}G$ barrelled by [18, 11.3.1 Proposition (e), p. 223]) if $F$ and $G$ are metrisable and barrelled by [18, 15.6.6 Proposition, p. 337], or if $F$ and $G$ are quasi-barrelled DF-spaces by [18, 15.6.8 Proposition, p. 338]. However, we emphasize that [15, Chap. I, §1, n1, Proposition 2, p. 30–31] in contrast to Theorem 4.5 does not need the restrictions that $E$ is complete and $F\otimes_{\pi}G$ quasi-barrelled. Moreover, Theorem 4.5 also complements [10, Theorem 3, p. 690] where a special continuous linearisation $(e,\mathcal{F}_{\ast}(\Omega),L)$ of a linear space $\mathcal{F}(\Omega)$ of scalar-valued continuous functions on a non-empty topological Hausdorff space $\Omega$ instead of a strong linearisation is considered and it is shown that the map $L(\mathcal{F}_{\ast}(\Omega),E)\to\mathcal{F}(\Omega,E)_{\sigma}\cap\mathcal{C}(\Omega,E)$, $u\mapsto u\circ e$, is an algebraic isomorphism if $E$ is a complete locally convex Hausdorff space. Here, $\mathcal{C}(\Omega,E)$ denotes the space of continuous maps from $\Omega$ to $E$. We refer the reader to [17, p. 182–184] for a summary of the construction of $(e,\mathcal{F}_{\ast}(\Omega),L)$
@@
If $Y$ is complete, then $Y$ is also a free object generated by $\Omega$ in the category $\mathcal{C}_{2}$ of complete locally convex Hausdorff spaces with continuous linear operators as morphisms in the sense of [12, p. 100] once the spaces $\Omega$ belong to a category $\mathcal{C}_{1}$ such that $\delta$ and any $f\in\mathcal{F}(\Omega,E)_{\sigma,b}$ is a morphism of $\mathcal{C}_{1}$ and there is a “forgetful” functor from $\mathcal{C}_{2}$ to $\mathcal{C}_{1}$.
@@
Let $(E,\tau)$ be a bornological locally convex Hausdorff space, $\mathcal{B}$ the family of $\tau$-bounded sets, $\widetilde{\tau}$ a locally convex Hausdorff topology on $E$ and $\widetilde{\gamma}$ denote the finest locally convex Hausdorff topology on $E$ which coincides with $\widetilde{\tau}$ on $\tau$-bounded sets. Suppose that a subset of $E$ is $\tau$-bounded if and only if it is $\widetilde{\gamma}$-bounded (see [25, Definition 3.12, p. 1600]). Then $(E,\widetilde{\gamma})^{\prime}$ determines boundedness of $(E,\tau)$. Indeed, $(E,\widetilde{\gamma})^{\prime}\subset(E,\tau)^{\prime}$ by [25, Remark 3.13 (c), p. 1600–1601]. Let $B\subset E$ be $\sigma(E,(E,\widetilde{\gamma})^{\prime})$-bounded. Then $B$ is $\widetilde{\gamma}$-bounded by Mackey’s theorem. It follows that $B$ is $\tau$-bounded by assumption.
@@
Let $(\mathcal{F}(\Omega),\tau)$ be a complete bornological DF-space of $\mathbb{K}$-valued functions on a non-empty set $\Omega$ satisfying $\operatorname{(BBC)}$ and $\operatorname{(CNC)}$ for some $\widetilde{\tau}$ such that $\tau_{\operatorname{p}}\leq\widetilde{\tau}$, $U\subset\Omega$ a set of uniqueness for $\mathcal{F}(\Omega)$, $E$ a complete locally convex Hausdorff space and $G\subset E^{\prime}$ determine boundedness. If $f\colon U\to E$ is a function such that $e^{\prime}\circ f$ admits an extension $f_{e^{\prime}}\in\mathcal{F}(\Omega)$ for each $e^{\prime}\in G$, then there exists a unique extension $F\in\mathcal{F}(\Omega,E)_{\sigma}$ of $f$.